Solve the compound inequality
Solve the compound inequality
Solve the compound inequality
Solve the absolute value equation
Solve the absolute value equation
Solve the absolute value equation
Solve the absolute value equation
Solve the inequality
Solve the inequality
Solve the inequality
Answer #1
1)
-20 + 5 5x 5 + 5
-155x10
-3x2
you can write this as [-3,2]
2)
x3 or 5x > 9 + 26
               5x>35
                 x> 7
you can write this as [3,) (7,)
3)
this is already solved. So, you can write this as (-1,3)
4)
31 - 4x = 18and          31-4x= -18
31-18 =4x                   31+ 18 = 4x  
13/4 =x                        49/4= x
3.25 =x                        12.25= x
So the solution are x = 3.25 and x = 12.25
5)
12x - 5 = 17 and       12x- 5 = -17
12x = 17 +5               12x= -17 + 5
x =22/12                        x= -12/12
                                       x= -1
So the solution are x = 22/12 and x = -1
6)
15n + 11 = 0
15n = -11
n = -11/15
7)
8v + 5 = -4 and       8v +5 = 4
8v =-4-5                 8v= 4 - 5
v =-9/8                     v= -1/8
So the solution are x = -9/8 and x = -1/8
8)
-13<2x-7<13
-13 + 7 < 2x < 13 + 7
-6<2x<20
-6/2
-3
9)
x + 6 - 13 or x + 6 13
x -13 + 6      x 13 - 6
x -7               x7
10)
20< 4x - 4 < -20
20 + 4 < 4x < -20 + 4
24<4x < -16
24/4 < x < -16/4
6